OpenStudy (anonymous):

4(7x - 1) = 3(2x +1) - 4(x + 6)

OpenStudy (anonymous):

keep in fraction and solve for x

OpenStudy (hoblos):

x= -17/26

OpenStudy (anonymous):

could you show me how you got that answer please?

OpenStudy (anonymous):

28x-4=6x+3-4x-24 28x-4=2x-21 26x=-17 x=-17/26

OpenStudy (hoblos):

=>28x-4=6x+3-4x-24 => 28x-4=2x-21 => 26x=-17 =>x=-17/26